OpenStudy (babyslapmafro):

If I am trying to find the area between two curves and on of the functions is x=(y^3)-y, how do I graph that function?

zepdrix (zepdrix):

I don't think there's a easy way to graph this. You certainly could take the derivative and find critical points and inflection points and graph it that way. But that sounds tedious. It might be a better idea to try and solve it without graphing it. It's possible you could be given a problem like that on a test - where it's not too hard to solve, but very very difficult to visualize. BTW, What is the other function? c:

OpenStudy (babyslapmafro):

x=0

OpenStudy (babyslapmafro):

maybe asking how to graph x=(y^3)-y is a bad question. How do I find out which is the rightmost bounding function?

zepdrix (zepdrix):

Hmm it turns out it's x=0... But I had to use a program to graph it to realize that. Hmmmmmmm... thinkingggg

zepdrix (zepdrix):

Oh ok let's try this.

zepdrix (zepdrix):

If we find intersecting points...\[\large x=y^3-y, \qquad x=0\]Setting these equal to one another will give us intersecting points.\[\large 0=y^3-y \qquad \rightarrow \qquad 0=y(y^2-1)\]So it looks like we have intersecting points at y=0, y=-1, y=1..... hmmm

zepdrix (zepdrix):

 Sorry I'm a little confused by this question... There appears to be 2 separate regions... see how one is to the left of x=0 and another to the right of x=0?